# Analytic Langlands correspondence from SoV

### v1

@@ 3.3 The inverse transformation
Unitarity of the $\mathsf{SoV}$-transformation ensures existence of an inverse transformation. It can be realised explicitly by inverting the relation (3.19), taking advantage of the fact that equations (3.18) can be used to define functions $y_{k}=y_{k}(\bm{k})$ up to permutations $y_{k}\mapsto y_{l}$, $y_{l}\mapsto y_{k}$. The resulting map $\mathsf{SoV}^{-1}$ takes elements $\phi$ of ${\mathcal{H}}_{\rm Skl}$ to scaling-invariant functions of n2n-2 variables $\bm{k}=(k_{2},\dots,k_{n-1})$.
@@
to eigenfunctions $\Psi_{\chi,i}$ of the differential operators $\mathsf{H}_{r}$, with eigenvalues $E_{r}$ determined by $t(y)$ using (2.9). One may furthermore easily check that $\phi_{\chi}(\bm{y})$ is normalizable, representing an element of ${\mathcal{H}}_{\rm Skl}$. This can be verified using (3.11), and taking into account that $\chi$ is analytic on $C$, having singular behaviour of the form
$\chi(y,\bar{y})=\sum_{\epsilon=\pm 1}|y-z_{r}|^{1+\epsilon(2j_{r}+1)}C_{r,\epsilon}(\bm{j})\big(1+{\mathcal{O}}(|y-z_{r}|\big),$ (3.33)
near the punctures $z_{r}$, $r=1,\dots,n$. Unitarity of the $\mathsf{SoV}$-transformation allows us to conclude that the eigenfunctions $\Psi_{\chi,i}=\mathsf{S}_{\bm{j},i}^{-1}(\phi_{\chi})$ are normalisable.
3.3.1 Smoothness
One should note, however, that the $L^{2}$-functions $\Psi_{\chi,i}=\mathsf{S}_{\bm{j},i}^{-1}(\phi_{\chi})$ are not necessarily point-wise defined, being defined by the Fourier-transformation of homogenous distributions.
@@ 3.3.1 Smoothness
$\mathsf{H}_{r}\Psi_{\chi,i}=E_{r}\Psi_{\chi,i},\quad r=2,\dots,n-2,\qquad\sum_{r=2}^{n-1}(x_{r}\partial_{r}-j_{r}x_{r})\Psi_{\chi,i}=0,$
@@
Elliptic regularity implies that $\Psi_{\chi,i}$ can be represented by functions $\Psi_{\chi,i}(\bm{x})$ which are smooth away from the singularities of the differential operators $\mathsf{H}_{r}$. Such singularities are associated to the so-called wobbly bundles [EFK3].
However, one should note that the type of singular behaviour that can occur at the wobbly locus is strongly constrained by the unitarity of the $\mathsf{SoV}$ transformation. As $\Psi_{\chi,i}$ are in the image of ${\mathcal{H}}_{\rm Skl}$ under the unitary operators $\mathsf{S}_{\bm{j},i}^{-1}$, it is ensured that the singularities of $\Psi_{\chi,i}$ can not spoil square-integrability. The unitarity of the $\mathsf{SoV}$-transformation allows us to avoid a detailed analysis of the singularities that may occur at the wobbly loci.
@@ 3.3.2 Meromorphic continuation
Let us finally note another property of the inverse transformation $\mathsf{S}_{\bm{j},i}^{-1}(\phi_{\chi})$ that will become important below. Considering the dependence of the wave-functions $\Psi_{\chi,i}$ with respect to the parameters $j_{1},\dots,j_{n}$, one may note that this dependence is meromorphic. In order to see this, one may first note that the dependence of $\chi$ on $j_{1},\dots,j_{n}$ is analytic, as $\chi$ is the solution to a system of differential equations having coefficients depending analytically on $j_{1},\dots,j_{n}$. The regular distribution associated to the function $\widetilde{\Psi}_{\bm{E}}(\bm{k})$ defined by inverting the relation (3.19) also depends analytically on $j_{1},\dots,j_{n}$. The inverse Fourier-transformation $\mathsf{F}_{\bm{j},i}^{-1}$ is meromorphic in $J$, as can be seen by an argument very similar to the one in [GrS, Section 2].
We may therefore consider the meromorphic continuation of wave-functions $\Psi_{\chi,i}$ away from the range $j_{r}\in-\frac{1}{2}+\mathrm{i}{\mathbb{R}}$ in which the unitarity of the $\mathsf{SoV}$-transformation holds.
@@ 4 Outline of the proof of the analytic Langlands correspondence
    (Separation of variables) Any eigenstate $\Psi_{\chi}\in{\mathcal{H}}_{i}$ of ${\mathsf{H}}_{w}$ defines an element $\Phi_{\chi}=\mathsf{S}_{\bm{j},i}\Psi_{\chi}$ of ${\mathcal{H}}_{\rm\mathsf{SoV}}$. The wave-function $\Phi_{\chi}(\bm{y})$ representing $\Phi_{\chi}$ has the form
@@
In order to prove these two results, we shall consider the $\mathsf{SoV}$ transformations $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}$, associated to $C_{0,n+1}=\mathbb{P}^{1}\setminus\{z_{0},\dots,z_{n}\}$, $\bm{j^{\prime}}=(j_{0},j_{1},\dots,j_{n})$. The definition of $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}$ is completely analogous to the definition of $\mathsf{S}_{\bm{j},i}^{(n)}:=\mathsf{S}_{\bm{j},i}$ given in the previous section, assigning index 0 to all variables associated to the additional puncture $z_{0}$, and admitting the factorisation $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}=\mathsf{SoV}^{(n+1)}\circ\mathsf{F}_{\bm{j},i}^{(n+1)}$. In the special case $j_{0}=-1$, we will exhibit relations between $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}$ and composition of $\mathsf{S}_{\bm{j},i}^{(n)}$ with the Hecke operators, implying the results i), ii) formulated above.
@@ 4.1 SoV-transform of eigenstates
The interplay between $\mathsf{SoV}$-transform and Hecke operators will prove one direction of the analytic Langlands correspondence rather easily.
@@
The $\mathsf{SoV}$-transform $\Phi_{\chi}=\mathsf{S}_{\bm{j},i}^{(n)}\Psi_{\chi}$ of an eigenstate $\Psi_{\chi}$ can be represented in terms of the eigenvalue $\chi(z_{0},\bar{z}_{0})$ of the Baxter operators $\mathsf{H}_{z_{0}}$ as
@@
can be used to define the $\mathsf{SoV}$ transform $\phi^{\prime}=\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}\Psi^{\prime}_{\chi}$ of the function $\Psi^{\prime}_{\chi}$ as before. The function $\phi^{\prime}(y_{0},\dots,y_{n-3})$ is manifestly symmetric in $y_{0},\dots,y_{n-3}$.
@@
As the $\mathsf{SoV}$-transform can be inverted, we see that eigenstates are uniquely determined by their eigenvalues, which means that the spectrum is simple.
@@ 4.2 Hecke eigenvalue property
We have seen that eigenstates define real opers as their Hecke eigenvalues. It remains to prove that all real opers define eigenstates. To this aim, we shall consider inverse $\mathsf{SoV}$ transformations of the form $\big(\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}\big)^{-1}\phi^{\prime}_{\chi}$, where
@@
$\big(\mathsf{S}_{\bm{j}^{\prime},0}^{(n+1)}\big)^{-1}\phi^{\prime}_{\chi}=\chi_{n}\,\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\,\mathbb{H}_{z_{0},x_{0}}\big(\mathsf{S}_{\bm{j},1}^{(n)}\big)^{-1}\phi_{\chi},$ (4.42)
using the notation $\chi_{n}:=C_{n,-}(\bm{j})$, referring to (3.33) for the definition of $C_{n,-}(\bm{j})$.
@@
The states $\psi_{\chi}=\big(\mathsf{S}_{\bm{j},i}^{(n)}\big)^{-1}\phi_{\chi}\in{\mathcal{H}}_{i}$ are eigenstates of the operators $\mathsf{Q}_{y}$,
$\mathsf{Q}_{w}\,\psi_{\chi}=\frac{1}{\chi_{n}}\chi(w,\bar{w})\,\psi_{\chi}.$ (4.43)
The proof will also be given in Section 6. We shall next turn to a few preparations.
5 An alternative integral representation
As a preparation for the proof of Theorem 3, we will here consider alternative representations for the eigenfunctions which do not use the simplifying choice $z_{n}=\infty$.
5.1 Definitions
A main ingredient of the integrands will be a single-valued function $\chi(y,\bar{y})$ satisfying
$\begin{aligned} &(\partial_{y}^{2}+t(y))\chi(y,\bar{y})=0,\\ &(\partial_{\bar{y}}^{2}+\bar{t}(\bar{y}))\chi(y,\bar{y})=0,\end{aligned}\qquad t(y)=\sum_{r=1}^{n}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{E_{r}}{y-z_{r}}\bigg).$ (5.44)

### v2

@@ 3.3 The inverse transformation
Unitarity of the SoV-transformation ensures existence of an inverse transformation. It can be realised explicitly by inverting the relation (3.20), taking advantage of the fact that equations (3.17) can be used to define functions $y_{k}=y_{k}(\bm{k})$ up to permutations $y_{k}\mapsto y_{l}$, $y_{l}\mapsto y_{k}$. The resulting map $\mathsf{SoV}^{-1}$ takes elements $\phi$ of ${\mathcal{H}}_{\rm Skl}$ to scaling-invariant functions $\Phi(\bm{k})$ of n2n-2 variables $\bm{k}=(k_{2},\dots,k_{n-1})$.
@@
to eigenfunctions $\Psi_{\chi,i}$ of the differential operators $\mathsf{H}_{r}$, with eigenvalues $E_{r}$ determined by $t(y)$ using (2.9). We shall next describe the main analytic properties of the functions $\Psi_{\chi,i}=\mathsf{S}_{\bm{j},i}^{-1}(\phi_{\chi})$.
3.3.1 Square-integrability
Square-integrability of the functions $\Psi_{\chi,i}=\mathsf{S}_{\bm{j},i}^{-1}(\phi_{\chi})$ will follow with the help of the unitarity of the SoV-transformation from the following lemma.
Lemma 3.
The functions $\phi_{\chi}$ satisfy
$\int_{{\mathbb{C}}^{n-3}}d\mu_{\rm Skl}(\bm{y})\;|\phi_{\chi}(\bm{y})|^{2}<\infty,$
and therefore represent elements of ${\mathcal{H}}_{\rm Skl}$.
Proof.
This can be verified using (3.11), and taking into account that $\chi$ is real analytic on $C$, having singular behaviour of the form
$\chi(y,\bar{y})=\sum_{\epsilon=\pm 1}|y-z_{r}|^{1+\epsilon(2j_{r}+1)}C_{\epsilon}^{(r)}(\bm{j})\big(1+{\mathcal{O}}(|y-z_{r}|)\big),$ (3.34)
near the punctures $z_{r}$, $r=1,\dots,n$. By using the explicit form of $d\mu_{\rm Skl}(\bm{y})$ given in (3.11) one may easily verify that $d\mu_{\rm Skl}(\bm{y})\;|\phi_{\chi}(\bm{y})|^{2}$ is integrable with respect to $y_{k}$ near $y_{k}=z_{r}$ for all $k=1,\dots,n-3$, and $r=1,\dots,n-1$, and furthermore
$d\mu_{\rm Skl}(\bm{y})\,|\phi_{\chi}(\bm{y})|^{2}=$ (3.35)
$\qquad=|y_{k}|^{-6}\cdot|y_{k}|^{4}\,d^{2}\bigg(\frac{1}{y_{k}}\bigg)\cdot\Bigg|\sum_{\epsilon=\pm 1}C_{\epsilon}^{(n)}(\bm{j})\,|y_{k}|^{1-\epsilon(2j_{n}+1)}\big(1+{\mathcal{O}}(y_{k}^{-1})\big)\Bigg|^{2}d\check{\mu}_{\rm Skl}(\check{\bm{y}}),$
with $\check{\bm{y}}=(y_{1},\dots,y_{k-1},y_{k+1},\dots,y_{n-3})$, and $d\check{\mu}_{\rm Skl}(\check{\bm{y}})$ obtained from $d\mu_{\rm Skl}(\bm{y})$ by omitting all factors containing $y_{k}$. Integrability with respect to $y_{k}$ near $z_{n}=\infty$ now follows from (3.35). ∎
Remark 1.
A coordinate-free description of $\mu_{\rm Skl}(\bm{y})$ has been proposed in [GT, Section 3.3.2].
3.3.2 Real-analyticity
One should note, however, that the $L^{2}$-functions $\Psi_{\chi,i}=\mathsf{S}_{\bm{j},i}^{-1}(\phi_{\chi})$ are not necessarily point-wise defined, being defined by the Fourier transformation of homogenous distributions.
@@ 3.3.2 Real-analyticity
$\mathsf{H}_{r}\Psi_{\chi,i}=E_{r}\Psi_{\chi,i},\quad r=2,\dots,n-2,\quad\sum_{r=1}^{n-1}(x_{r}\partial_{x_{r}}-j_{r})\Psi_{\chi,i}=0,\quad\sum_{r=1}^{n-1}\partial_{x_{r}}\Psi_{\chi,i}=0,$
@@
Elliptic regularity for holonomic systems implies that $\Psi_{\chi,i}$ can be represented by functions $\Psi_{\chi,i}(\bm{x})$ which are real-analytic away from the singular locus of the differential operators $\mathsf{H}_{r}$.
Remark 2.
The points in the singular locus correspond to the so-called wobbly bundles [EFK3]. However, one should note that the type of singular behaviour that can occur at the wobbly locus is strongly constrained by the unitarity of the SoV transformation. As $\Psi_{\chi,i}$ are in the image of ${\mathcal{H}}_{\rm Skl}$ under the unitary operators $\mathsf{S}_{\bm{j},i}^{-1}$, it is ensured that the singularities of $\Psi_{\chi,i}$ can not spoil square-integrability. The unitarity of the SoV-transformation allows us to avoid a detailed analysis of the singularities that may occur at the wobbly loci.
@@ 3.3.3 Meromorphic continuation
Let us finally note another property of the inverse transformation $\mathsf{S}_{\bm{j},i}^{-1}(\phi_{\chi})$ that will become important below. Considering the dependence of the wave-functions $\Psi_{\chi,i}$ with respect to the parameters $j_{1},\dots,j_{n}$, one may note that this dependence is meromorphic. In order to see this, one may first note that the dependence of $\chi$ on $j_{1},\dots,j_{n}$ is analytic, as $\chi$ is the solution to a system of differential equations having coefficients depending analytically on $j_{1},\dots,j_{n}$. The regular distribution associated to the function $\widetilde{\Psi}_{\bm{E}}(\bm{k})$ defined by inverting the relation (3.20) also depends analytically on $j_{1},\dots,j_{n}$. The inverse Fourier transformation $\mathsf{F}_{\bm{j},i}^{-1}$ is meromorphic in $J$, as can be seen by an argument very similar to the one in [GrS, Section 2].
We may therefore consider the meromorphic continuation of wave-functions $\Psi_{\chi,i}$ away from the range $j_{r}\in-\frac{1}{2}+\mathrm{i}{\mathbb{R}}$ in which the unitarity of the SoV-transformation holds.
@@ 4 Outline of the proof of the analytic Langlands correspondence
    (Separation of variables) Any eigenstate $\Psi_{\chi}\in{\mathcal{H}}_{i}$ of ${\mathsf{H}}_{w}$ defines an element $\phi_{\chi}=\mathsf{S}_{\bm{j},i}\Psi_{\chi}$ of ${\mathcal{H}}_{\rm SoV}$. The wave-function $\phi_{\chi}(\bm{y})$ representing $\phi_{\chi}$ has the form
@@
In order to prove these two results, we shall consider the SoV transformations $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}$, associated to $C_{0,n+1}=\mathbb{P}^{1}\setminus\{z_{0},\dots,z_{n}\}$, $\bm{j^{\prime}}=(j_{0},j_{1},\dots,j_{n})$. The definition of $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}$ is completely analogous to the definition of $\mathsf{S}_{\bm{j},i}^{(n)}:=\mathsf{S}_{\bm{j},i}$ given in the previous section, assigning index 0 to all variables associated to the additional puncture $z_{0}$, and admitting the factorisation $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}=\mathsf{SoV}^{(n+1)}\circ\mathsf{F}_{\bm{j},i}^{(n+1)}$. In the special case $j_{0}=-1$, we will exhibit relations between $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}$ and the composition of $\mathsf{S}_{\bm{j},i}^{(n)}$ with the Hecke operators, implying the results i), ii) formulated above.
@@ 4.1 SoV-transform of eigenstates
The interplay between SoV-transform and Hecke operators will prove one direction of the analytic Langlands correspondence rather easily.
@@
The SoV-transform $\phi_{\chi}=\mathsf{S}_{\bm{j},i}^{(n)}\Psi_{\chi}$ of an eigenstate $\Psi_{\chi}$ can be represented in terms of the eigenvalue $\chi(z_{0},\bar{z}_{0})$ of the Baxter operators $\mathsf{H}_{z_{0}}$ as
@@
can be used to define the SoV transform $\phi^{\prime}=\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}\Psi^{\prime}_{\chi}$ of the function $\Psi^{\prime}_{\chi}$ as before. The function $\phi^{\prime}(y_{0},\dots,y_{n-3})$ is manifestly symmetric in $y_{0},\dots,y_{n-3}$.
@@
As the SoV-transform can be inverted, we see that eigenstates are uniquely determined by their eigenvalues, which means that the spectrum is simple.
@@ 4.2 Hecke eigenvalue property
We have seen that eigenstates define real opers as their Hecke eigenvalues. It remains to prove that all real opers define eigenstates. To this aim, we shall consider inverse SoV transformations of the form $\big(\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}\big)^{-1}\phi^{\prime}_{\chi}$, where
@@
$\mathbb{H}_{z_{0},x_{0}}\big(\mathsf{S}_{\bm{j},0}^{(n)}\big)^{-1}\phi_{\chi}=\frac{1}{\chi_{n}}\big(\mathsf{S}_{\bm{j}^{\prime},1}^{(n+1)}\big)^{-1}\phi^{\prime}_{\chi},\qquad\chi_{n}:=\frac{\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\,C_{-}^{(n)}(\bm{j}),$ (4.44)
referring to (3.34) for the definition of $C_{-}^{(n)}(\bm{j})$.
@@
The states $\psi_{\chi}=\big(\mathsf{S}_{\bm{j},i}^{(n)}\big)^{-1}\phi_{\chi}\in{\mathcal{H}}_{1}$ are eigenstates of the operators $\mathbf{Q}_{w}$,
$\mathbf{Q}_{w}\,\psi_{\chi}=\frac{1}{\chi_{n}}\chi(w,\bar{w})\,\psi_{\chi}.$ (4.45)
The proof will be given in Section 6.
5 Hecke modifications of the inverse SoV transformation
We shall here prove the crucial Theorem 3. We will consider the inverse $\mathsf{SoV}$-transformations associated to Riemann spheres $C_{0,n+1}$ obtained from $C_{0,n}$ by adding another marked point at $z=z_{0}$. When the parameter $j_{0}$ associated to $z_{0}$ is chosen as $j_{0}=-1$, we will find a relation with the Hecke modification $\mathbb{H}_{z_{0},x_{0}}$.
We begin by introducing the relevant notations. A main ingredient of the integrand will be a single-valued function $\chi(y,\bar{y})$ satisfying
$\begin{aligned} &(\partial_{y}^{2}+t(y))\chi(y,\bar{y})=0,\\ &(\partial_{\bar{y}}^{2}+\bar{t}(\bar{y}))\chi(y,\bar{y})=0,\end{aligned}\qquad t(y)=\sum_{r=1}^{n-1}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{E_{r}}{y-z_{r}}\bigg).$ (5.46)